# Symbolic Mixture-of-Experts: Adaptive Skill-based Routing for Heterogeneous Reasoning

### v2

@@ Abstract
Combining existing pre-trained expert LLMs is a promising avenue for scalably tackling large-scale and diverse tasks. However, selecting experts at the task level is often too coarse-grained, as heterogeneous tasks may require different expertise for each instance. To enable adaptive instance-level mixing of pre-trained LLM experts, we propose Symbolic-MoE, a symbolic, text-based, and gradient-free Mixture-of-Experts framework. Symbolic-MoE takes a fine-grained approach to selection by emphasizing skills, i.e., specialized subcategories or subtopics such as algebra in mathematics or molecular biology in biomedical reasoning. We propose a skill-based recruiting strategy that dynamically selects the most relevant set of expert LLMs for diverse reasoning tasks based on their strengths. Each selected expert then generates its own reasoning, resulting in $k$ outputs from $k$ experts, which are then synthesized into a final high-quality response by an aggregator. The aggregator is chosen based on its ability to integrate diverse reasoning outputs. We show that instance-level expert selection improves performance by a large margin but -- when implemented naively -- can introduce a high computational overhead due to the need for constant model loading and offloading. To address this, we implement a batch inference strategy that groups instances based on their assigned experts, ensuring each model will only be loaded once. This allows us to integrate 16 models on a single GPU with a time cost comparable to prior multi-agent baselines using 4 GPUs. Through extensive evaluations on diverse benchmarks (MMLU-Pro, GPQA, AIME, and MedMCQA), we demonstrate that Symbolic-MoE outperforms strong LLMs like GPT4o-mini, as well as multi-agent approaches, with an absolute average improvement of $8.15\%$ over the best multi-agent baseline. Moreover, Symbolic-MoE removes the need for expensive multi-round discussions, outperforming discussion baselines with less computation.
@@ 1 Introduction
A core strength of humans is our ability to communicate and coordinate with each other using language (Clark, 1996; Yow & Lim, 2019; Xu et al., 2023). This allows diverse experts to contribute specialized knowledge towards solving a problem, and is common across a variety of settings, including research, medicine, and engineering. Like humans, large language models (LLMs) often have differing skills and strengths, derived from differences in their architectures and training regimens. For instance, math-specific models like MetaMath (Yu et al., 2023), WizardMath (Luo et al., 2023), and QwenMath (Yang et al., 2024) are post-trained with mathematical reasoning data, making them particularly adept at math tasks -- often at the cost of performance on out-of-distribution tasks (Kumar et al., 2022; Chu et al., 2025) like commonsense or medical reasoning (Lobo et al., 2024). Even within specialized domains, differences in pre-training data can lead to nuanced variations in expertise: one math-focused model may excel at algebra, while another is better suited for geometry. This motivates our development of an automated, skill-based framework designed to identify and select the most suitable set of expert models for a given problem. [^1] We refer experts to those LLMs that are selected to solve a problem in this work. Allowing multiple diverse models to combine their answers can improve even the strongest models (Chen et al., 2024c), and as tasks grow in complexity, leveraging specialized models at test time presents a promising approach to scaling LLM capabilities.
@@
Indeed, combining multiple ‘‘expert’’ models, i.e., Mixture-of-Experts (MoE) is a well-studied notion in machine learning (Jacobs et al., 1991; Eigen et al., 2013) and has been applied widely for large pre-trained models, enabling better performance at a lower computing cost (Shazeer et al., 2017a; Fedus et al., 2022; Riquelme et al., 2021). However, in the conventional MoE settings, experts are typically sub-models, i.e. subsets of parameters within a larger model, trained simultaneously in an end-to-end fashion, where at test time, experts are combined in the model’s parameter space. This generally requires end-to-end training from scratch, which is often computationally expensive and precludes the re-use of the vast pool of existing already-trained LLMs. Building on recent efforts in combining a fixed set of models through multi-agent discussions (Chen et al., 2024c; Du et al., 2023; Liang et al., 2023; Wang et al., 2024a), we propose exploring a new training-free paradigm for large-scale MoEs: a symbolic mixture of experts (Symbolic-MoE). Rather than using information encoded in the model’s hidden state, Symbolic-MoE uses symbolic structures in two ways: First, Symbolic-MoE infers a set of discrete skills needed to solve a problem, measuring the abilities of each model in a pool of candidate expert models according to the set. It then uses skill-based performance as a ‘‘router’’ to recruit a sparse subset of experts for each problem. Secondly, Symbolic-MoE combines pretrained experts through a symbolic channel, i.e., language, which is a common protocol already shared by all models involved. To take advantage of the diverse set of expert LLMs, two key challenges present: (1) Effective Expert Selection: Given a large set of existing LLMs, how can we choose the best experts for each example? (2) Scalable Expert Mixing: How can we scalably serve a large number of experts without increasing the demand for GPUs?
@@
Effective Expert Selection: The increasing diversity of benchmarks (Miranda et al., 2024) and the growing number of models means that experts must be selected not at the level of tasks, but at the level of individual queries. Even at the task level, manual selection can be labor-intensive, and the performance of multi-agent frameworks can be sensitive to the agents chosen for a task (Chen et al., 2024c; Liang et al., 2023; Wang et al., 2024b). At the query level, manual selection becomes infeasible, especially in multi-model settings where different experts must collaborate to address complex queries (Wang et al., 2024c; Rein et al., 2023). Moreover, manually selecting models for each task does not guarantee optimal performance. For instance, as shown in Fig. 1 (a), while a given subset of models may perform well on math tasks on average, their proficiency in specific subfields like algebra or probability might vary. Therefore, using of a fixed subset of models on all math samples might hurt performance on particular subtasks. This underscores the need for an automated, fine-grained selection mechanism as shown in Fig. 1 (b). Given that evaluating all possible model subsets is computationally infeasible, we instead propose an efficient approach that activates only the most relevant experts while minimizing computational overhead.
@@
Figure 1: Comparison between existing multi-agent work and our work. (a) In prior work, a fixed set of task-level experts (e.g., Phi, Mistral, and Llama) is recruited to solve mathematical problems, where heterogeneous questions may differ in the skills required to solve them (e.g Q1 requires algebra, while Q2 focuses on probability). Expert models then engage in multiple rounds of discussion, making this approach resource-intensive. (b) In contrast, Symbolic-MoE adaptively recruits instance-level experts based on skills needed (‘‘Algebra" experts for Q1 and a different set of ‘‘Probability" experts for Q2). By generating only a single round of responses with a selected aggregator to synthesize the final output, our approach is both more performant and more efficient.
Scalable Expert Mixing: Unless models are retrained jointly, the integration of multiple models must occur through their outputs -- typically via a symbolic channel, such as multi-agent discussion. Past work has often relied on multiple rounds of inference, leading to significant GPU demands. As shown in Fig. 3 (I), when selecting a fixed set of models for a given domain (e.g. as done by (Li et al., 2024b; 2025)), models could in principle be loaded onto parallel GPUs. This reduces the latency of the system but increases the number of GPUs needed. Moreover, it does not scale to a dynamic setting like the one we consider, where the number of GPUs required would be equal to the number of potential models available (in our case, $16$), making this option prohibitively expensive. Given $k$ models and $n$ queries, this solution requires $k$ GPUs. Alternatively, models could be run sequentially on a single GPU (cf. Fig. 3 (II)), which minimizes GPU usage but incurs high overhead due to frequent loading and off-loading of models. In other words, this solution requires $\mathcal{O}(kn)$ loads. Given the high cost of loading models into GPU memory (Griggs et al., 2024; Li et al., 2024a), this solution is also untenable. Processing each query separately also fails to fully utilize GPU memory, further limiting its efficiency.
Driven by these motivations, in Symbolic-MoE, we achieve effective expert selection and scalable expert mixing through the following approach. First, to enhance expert selection, we introduce automated skill-based recruiting, which leverages an inferred skill-based model profile -- a dictionary mapping the most relevant models to specific skills (e.g., keywords or subcategories) required to solve a given query -- using a small subset of available samples. Like a router in a standard MoE setting, Symbolic-MoE’s skill-based routing allows the most relevant $k$ experts to be activated while keeping the others inactive, optimizing both accuracy and efficiency. Next, to enable scalable expert mixing, we propose a batch inference mechanism (illustrated in Fig. 3 (III)). For each query, the router first selects the necessary experts, and then examples are grouped into batches per model. We then run all queries for a given model in a single batch, allowing us to load and off-load $\mathcal{O}(k)$ times in total, which is far faster than sequential processing’s $\mathcal{O}(kn)$. Finally, while Symbolic-MoE’s batched implementation accommodates up to $16$ models on a single GPU, it can also be parallelized across multiple GPUs. This flexibility ensures both speedups with increased computing power and accessibility for users with limited hardware resources. As shown in Figure 1, existing multi-agent approaches use a fixed set of experts for each task (e.g., all math problems), regardless of fine-grained topics (e.g., algebra or probability), and rely on multiple discussion rounds, making them non-adaptive and resource-intensive. In contrast, our approach introduces adaptive query-level recruitment based on fine-grained skills, selecting the most suitable models for each query. Furthermore, we improve efficiency through (1) a batch inference strategy that processes all queries for a model in a single batch and (2) a task-specific aggregator instead of multi-agent discussions, ensuring that expert models generate their answers only once.
@@
We evaluate Symbolic-MoE on diverse benchmarks, including MMLU-Pro (Wang et al., 2024c), GPQA (Rein et al., 2023), AIME 2024 (MAA, 2024), and MedMCQA Pal et al. (2022), using a diverse model pool. We show that integrating this model pool with automated skill-based recruiting yields an average accuracy improvement of $8.15\%$ over the best multi-agent baseline. Moreover, despite primarily using LLMs with 7-8 billion (B) parameters, Symbolic-MoE achieves comparable performance with larger 70B models, and on average, outperforms strong proprietary models like GPT4o-mini (OpenAI, 2024), Gemini 1.5 Pro (Team et al., 2024), and DeepSeek-V3 (DeepSeek-AI et al., 2025b). Without any manual intervention, Symbolic-MoE consistently surpasses all baselines, whereas the strongest baseline changes across settings, with some baselines performing competitively in one setting but poorly in others. Thus, Symbolic-MoE eliminates the need for the user to evaluate and compare a large number of possible baselines for each setting. Notably, Symbolic-MoE obtains these benefits while reducing the amount of compute needed. Using a single GPU, Symbolic-MoE has $44\%$ less run-time than a mixture-of-agents baseline (Wang et al., 2024a). When four GPUs are available for both methods, we obtain an almost 2×2$\times$2 × speedup over this baseline. Our analysis highlights that Symbolic-MoE shows strong robustness despite the variation in the best-performing models and optimal aggregators for each task. Additionally, we show that selecting a task-specific aggregator based on its ability to integrate diverse answers can achieve performance comparable to multi-round discussions while requiring substantially less compute.
@@ 2 Related Work
Traditional Mixture-of-Experts (MoE) (Jacobs et al., 1991; Jordan & Jacobs, 1994; Chen et al., 1999; Yuksel et al., 2012) models distribute computation across multiple experts, with growing interest in sparsity-driven approaches. The Sparse MoE (SMoE) approach (Shazeer et al., 2017a) improves efficiency by activating only the most relevant experts per input, enhancing scalability for high-dimensional data. This method has been widely applied in vision tasks (Riquelme et al., 2021; Wang et al., 2020; Yang et al., 2019; Abbas & Andreopoulos, 2020), language tasks (Lepikhin et al., 2021; Zhang et al., 2021; Zuo et al., 2022; Jiang et al., 2021) and multimodal learning (Kudugunta et al., 2021; Yun et al., 2024). In contrast to past sparse MoE work, the models in our approach are combined through symbolic channels (e.g., language outputs) as opposed to in the parameter space, allowing us to re-use existing LLMs without training. MoA (Wang et al., 2024a) introduces a framework for combining LLM agents into ensembles that relies on a fixed set of agents across tasks and instances. This approach requires multiple rounds of generation and aggregation before producing a final answer. Similarly, Self-MoA  (SMoA; Wang et al., 2024a) posits that using multiple distinct LLMs is unnecessary, suggesting that optimal performance can be achieved by invoking the task-best model multiple times alongside the task-best aggregator. Our work differs from MoA and SMoA by introducing adaptive, instance-level, skill-based routing while avoiding costly multi-model discussions in favor of streamlined aggregation, significantly reducing computational overhead. We also find that mixing different LLMs is advantageous when paired with effective routing and aggregation strategies. Moreover, we show that the best aggregator for a task is not necessarily the best-performing model overall, but can be identified through a synthetic task we introduce, designed to evaluate aggregation effectiveness.
Multi-Agent Reasoning.
@@
Multi-agent reasoning has emerged as a promising paradigm for enhancing complex problem-solving and decision-making in AI systems. Early approaches employed reinforcement learning-based coordination (Lowe et al., 2017; Foerster et al., 2018; Jaques et al., 2019), while recent efforts leverage LLM-based multi-agent frameworks. One line of research explores student-teacher paradigms (Magister et al., 2022; Fu et al., 2023; Ho et al., 2022; Du et al., 2023; Chen et al., 2024a), where reasoning capabilities are distilled from stronger to weaker agents. Another approach investigates multi-agent debate frameworks, where agents interact to refine arguments and enhance collective decision-making; this has been explored with multiple instances of a single model (Liang et al., 2023; Xiong et al., 2023; Chan et al., 2023) or debates between multiple LLM types (Chen et al., 2024c). In both cases, the set of models is predefined by the user. In contrast, our approach automatically selects models based on inferred skills. Additionally, our framework achieves superior efficiency by avoiding multi-round discussions while still outperforming debate-based methods.
@@ 3 Methodology
Symbolic-MoE consists of three stages: (I) model profile creation and aggregator selection (Fig. 2 (a)), which occur during preprocessing, followed by (II) expert recruitment and (III) final answer generation (Fig. 2 (b)), both of which take place during inference. First, we establish the problem setup in §$\S$3.1. Then, in the preprocessing stage (§$\S$3.2), we describe how the model profile is created in §$\S$3.2.1, which enables a fine-grained skill-based assessment of each model. We also describe the process for selecting the aggregator in §$\S$3.2.2, which is responsible for generating the final answer. Next, in the inference stage (§$\S$3.3), we introduce the recruitment process in §$\S$3.3.1, where the experts are selected based on the skills needed for each problem, and how the answers from the recruited experts are combined using the aggregator to improve reasoning in §$\S$3.3.2.
@@ 3.1 Problem Setup
Given a pool of $n$ models =$\mathcal{M}=\{M_{i}\}_{i=1}^{n}$, where each model represents a distinct LLM with potentially different pre-training datasets and architectures, our goal is to optimize performance through dynamic allocation -- solving each problem with the most suitable subset of $k$ models -- and combined reasoning -- allowing experts to combine information to enhance reasoning. To achieve this, we assume access to a small validation set to obtain (1) model profiles $P_{i}\;\forall i\leq n$, and (2) aggregator profiles that benchmark the ability of each model to act as an aggregator. We use these profiles to recruit experts (at the instance level) and to select the aggregator (at the task level).

### v3

@@ Abstract
Combining existing pre-trained expert LLMs is a promising avenue for scalably tackling large-scale and diverse tasks. However, selecting experts at the task level is often too coarse-grained, as heterogeneous tasks may require different expertise for each instance. To enable adaptive instance-level mixing of pre-trained LLM experts, we propose Symbolic-MoE, a symbolic, text-based, and gradient-free Mixture-of-Experts framework. Symbolic-MoE takes a fine-grained approach to selection by emphasizing skills, i.e., specialized subcategories such as algebra in mathematics. We propose a skill-based recruiting strategy that dynamically selects the most relevant set of expert LLMs for diverse reasoning tasks based on their strengths. Each selected expert then generates its own reasoning, resulting in $k$ outputs from $k$ experts, which are then synthesized into a final high-quality response by an aggregator, which is chosen based on its ability to integrate diverse outputs. We show that instance-level expert selection improves performance by a large margin but – when implemented naively – can introduce a high computational overhead due to the need for constant model loading and offloading. To address this, we implement a batch inference strategy that groups instances based on their assigned experts, ensuring each model will only be loaded once. This allows us to integrate 16 models on a single GPU with a time cost comparable to prior multi-agent baselines using 4 GPUs. Through extensive evaluations on diverse benchmarks (MMLU-Pro, GPQA, AIME, and MedMCQA), we show that Symbolic-MoE outperforms prior multi-agent approaches, with an absolute average improvement of $8.15\%$ over the best baseline. Moreover, Symbolic-MoE generalizes well to unseen tasks and removes the need for expensive multi-round discussions, outperforming discussion baselines with less computation. [^1] Project page: https://symbolic-moe.github.io
@@ 1 Introduction
A core strength of humans is our ability to communicate and coordinate with each other using language (Clark, 1996; Yow & Lim, 2019; Xu et al., 2023). This allows diverse experts to contribute specialized knowledge towards solving a problem. Like humans, large language models (LLMs) often have differing skills and strengths, derived from differences in their architectures and training regimens. For instance, math-specific models like MetaMath (Yu et al., 2023) or QwenMath (Yang et al., 2024) are post-trained with mathematical reasoning data, making them particularly adept at math tasks – often at the cost of performance on out-of-distribution tasks (Kumar et al., 2022; Chu et al., 2025) like commonsense or medical reasoning (Lobo et al., 2024). Even within specialized domains, differences in pre-training data can lead to nuanced variations in expertise: one math-focused model may excel at algebra, while another is better suited for geometry. This motivates our development of an automated, skill-based framework designed to identify and select the most suitable set of expert models for each problem.
@@
Indeed, combining multiple “expert” models via Mixture-of-Experts (MoE) is well-studied (Jacobs et al., 1991; Eigen et al., 2013) and has been applied widely for large pre-trained models, enabling better performance at a lower computational cost (Shazeer et al., 2017a; Fedus et al., 2022; Riquelme et al., 2021). However, in the conventional MoE settings, experts are typically sub-models, i.e. subsets of parameters within a larger model, where at test time, they are combined in the model’s parameter space. This generally requires end-to-end training from scratch, which is often computationally expensive and precludes the re-use of the vast pool of existing already-trained LLMs. Building on recent efforts in combining a fixed set of models through multi-agent discussions (Chen et al., 2024c; Du et al., 2023; Liang et al., 2023; Wang et al., 2024a), we propose exploring a new training-free paradigm for large-scale MoEs: a symbolic mixture of experts (Symbolic-MoE). Rather than using information encoded in the model’s hidden state, Symbolic-MoE uses symbolic structures in two ways: First, Symbolic-MoE infers a set of discrete skills needed to solve a problem, measuring the abilities of each model in a pool of candidate expert models. It then uses skill-based performance as a “router” to recruit a sparse subset of experts for each problem. Secondly, Symbolic-MoE combines pre-trained experts through a symbolic channel, i.e., language, which is a common protocol already shared by all LLMs. To take advantage of the diverse set of expert LLMs, we must address two key: (1) Effective Expert Selection: Given a large set of existing LLMs, how can we choose the best experts for each instance? (2) Scalable Expert Mixing: How can we serve a large number of experts (e.g. $16$) without increasing the demand for GPUs?
@@
(1) Effective Expert Selection: The increasing diversity of benchmarks (Miranda et al., 2024) and the growing number of models means that experts must be selected not at the level of tasks, but at the level of individual queries. Even at the task level, manual selection can be labor-intensive, and the performance of multi-agent frameworks can be sensitive to the agents chosen for a task (Chen et al., 2024c; Liang et al., 2023; Wang et al., 2024b). For instance, as shown in Fig. 1 (a), while a given subset of models may perform well on math tasks on average, their proficiency in specific subfields like algebra or probability might vary – that is, using a fixed subset of models on all math samples might hurt performance on particular subtasks. This underscores the need for an automated, fine-grained selection mechanism, as shown in Fig. 1 (b).
@@
Figure 1: (a) In prior work, a fixed set of task-level experts is recruited to solve mathematical problems, while heterogeneous questions may differ in the skills required to solve them (e.g., Q1 requires algebra, while Q2 focuses on probability). The recruited experts generate outputs for multiple rounds, making these methods inefficient. (b) In contrast, Symbolic-MoE adaptively recruits instance-level experts based on skills needed (“Algebra” experts for Q1 and a different set of “Probability” experts for Q2). By generating only a single round of responses with an aggregator to synthesize the final output, our approach is both more performant and more efficient.
(2) Scalable Expert Mixing: Past work has often relied on multiple rounds of inference, leading to significant GPU demands. Moreover, it does not scale to a dynamic setting like the one we consider, where the number of GPUs required would be equal to the number of potential models available (in our case, $16$), making this option prohibitively expensive. We instead propose a batch inference mechanism (see Fig. 3 (III)). Specifically, the samples are grouped into batches per model. We then run all queries assigned to a given model in a single batch, which is far faster than sequential processing. While this strategy accommodates up to $16$ models on a single GPU, it can also be parallelized across multiple GPUs. This flexibility ensures both speedups with increased computing power, and accessibility for users with limited hardware resources.
@@
We evaluate Symbolic-MoE on diverse benchmarks, including MMLU-Pro (Wang et al., 2024c), GPQA (Rein et al., 2023), AIME 2024 (MAA, 2024), and MedMCQA Pal et al. (2022), using a diverse model pool. We show that our automated skill-based recruiting yields an average accuracy improvement of $8.15\%$ over the best multi-agent baseline. Moreover, despite primarily using LLMs with 7-8 billion (B) parameters, Symbolic-MoE achieves comparable performance with 70B models, and on average, outperforms strong proprietary models like GPT4o-mini (OpenAI, 2024), Gemini 1.5 Pro (Team et al., 2024a), and DeepSeek-V3 (DeepSeek-AI et al., 2025b). Also, Symbolic-MoE consistently surpasses all baselines, whereas the strongest baseline changes across tasks. Thus, Symbolic-MoE eliminates the need for the user to evaluate and compare a large number of possible baselines for each task. Notably, using a single GPU, Symbolic-MoE has $44\%$ less inference run-time than a mixture-of-agents baseline (Wang et al., 2024a); when four GPUs are available for both methods, we obtain an almost 2×2$\times$2 × speedup. Finally, our analysis shows that Symbolic-MoE generalizes well to unseen tasks, and selecting a task-specific aggregator achieves performance comparable to multi-round discussion while requiring substantially less compute.
@@ 2 Related Work
Mixture-of-Agents. Traditional Mixture-of-Experts (MoE) models (Jacobs et al., 1991; Jordan & Jacobs, 1994; Chen et al., 1999; Yuksel et al., 2012) distribute computation across multiple “expert” submodels, with growing interest in sparsity-driven approaches. The Sparse MoE (SMoE) approach (Shazeer et al., 2017a) improves efficiency by activating only the most relevant experts per input, enhancing scalability for high-dimensional data. This method has been widely applied in vision tasks (Riquelme et al., 2021; Wang et al., 2020; Yang et al., 2019; Abbas & Andreopoulos, 2020), language tasks (Lepikhin et al., 2021; Zhang et al., 2021; Zuo et al., 2022; Jiang et al., 2021) and multimodal learning (Kudugunta et al., 2021; Yun et al., 2024). Unlike Symbolic-MoE, standard MoE approaches require experts to be trained jointly, with communication taking place in the parameter spaces of the submodels. On the test-time mixture side, MoA (Wang et al., 2024a) combine LLM agents into ensembles that rely on a fixed set of agents. This approach requires multiple rounds of generation and aggregation before producing a final answer. Similarly, Self-MoA  (Li et al., 2025) suggest that optimal performance can be achieved by invoking the task-best model multiple times alongside the task-best aggregator. Our work differs from MoA and Self-MoA by introducing adaptive, instance-level, skill-based routing while avoiding costly multi-model discussions in favor of streamlined aggregation. We also find that mixing different LLMs is advantageous when paired with effective routing and aggregation strategies. Moreover, we show that the best aggregator for a task is not necessarily the best-performing model overall, but can be identified through a synthetic task we introduce, designed to evaluate aggregation effectiveness.
@@
Multi-Agent Reasoning. Multi-agent reasoning has emerged as a promising paradigm for enhancing complex problem-solving and decision-making in AI systems. Early approaches employed reinforcement learning-based coordination (Lowe et al., 2017; Foerster et al., 2018; Jaques et al., 2019), while recent efforts leverage LLM-based multi-agent frameworks. One line of research explores student-teacher paradigms (Magister et al., 2022; Fu et al., 2023; Ho et al., 2022; Du et al., 2023; Chen et al., 2024a), where reasoning capabilities are distilled from stronger to weaker agents. Another approach investigates multi-agent debate frameworks, where agents interact to refine arguments and enhance collective decision-making; this has been explored with multiple instances of a single model (Liang et al., 2023; Xiong et al., 2023; Chan et al., 2023) or debates between multiple LLM types (Chen et al., 2024c). In both cases, the set of models is predefined by the user. In contrast, our approach automatically selects models based on inferred skills. Additionally, our framework achieves superior efficiency by avoiding multi-round discussions while still outperforming debate-based methods.
@@ 3.1 Problem Setup
Given a pool of $n$ models =$\mathcal{M}=\{M_{i}\}_{i=1}^{n}$, where each model represents a distinct LLM with potentially different pre-training datasets and architectures, our goal is to optimize performance through dynamic allocation – solving each problem with the most suitable subset of $k$ models, allowing experts to combine information to enhance reasoning. To achieve this, we assume access to a small validation set to obtain (1) model profiles $P_{i}\;\forall i\leq n$, and (2) aggregator profiles that benchmark the ability of each model to act as an aggregator. We use these profiles to recruit experts (at the instance level) and to select the aggregator (at the task level).